OpenStudy (anonymous):

use the quadratic formula to solve equation 5x^2-7x-3=0 what is the slution set

OpenStudy (anonymous):

\[ \left\{\left\{x= \frac{1}{10} \left(7-\sqrt{109}\right)\right\},\left\{x= \frac{1}{10} \left(7+\sqrt{109}\right)\right\}\right\} \]